# Vid-LLM: A Compact Video-based 3D Multimodal LLM with Reconstruction–Reasoning Synergy

### v1

@@ Appendix A Appendix
  • qualitative visualizations of Vid-LLM outputs in Appendix A.1;
@@
  • comprehensive implementation details, covering optimization settings, training procedures, and hyperparameters in Appendix A.3.
@@
A.1 Qualitative Results
@@
Fig. 7 presents reconstruction results on the ScanNet dataset, together with three representative 3D vision–language (3D VL) tasks. Specifically, it illustrates (1) 3D Question Answering, where the model answers queries on object counts, locations, and attributes from reconstructed scenes; (2) 3D Dense Captioning, which provides detailed semantic descriptions of rooms and key objects; and (3) 3D Visual Grounding, where the model localizes target objects in 3D space according to textual instructions. These qualitative results demonstrate the model’s capability in scene reconstruction and 3D VL reasoning from video inputs.
A.2 Supplementary Experiments on Reconstruction
Overview. Since the reconstruction branch not only provides 3D cues for the 3D-MLLM but also explicitly produces point clouds that can serve downstream applications, it is important to assess its individual performance. To evaluate the 3D modeling capability of Vid-LLM, we assess three tasks: camera pose estimation, depth prediction, and point-cloud reconstruction. Depth prediction is crucial to real-scale reconstruction, and together with poses it determines the quality of the recovered point clouds. Following standard protocols, we evaluate poses on Co3Dv2 (Reizenstein et al., 2021) and RealEstate10K (Zhou et al., 2018), depth on the indoor NYU Depth v2 dataset (Silberman et al., 2012), and point-cloud reconstruction on ScanNet (Dai et al., 2017), ensuring consistency with the 3D vision–language benchmarks introduced earlier.
Baseline. For 3D reconstruction, we evaluate three subtasks: (i) camera pose estimation, where we compare against geometry-based optimization methods including COLMAP+SG and PixSfM, as well as state-of-the-art learning-based approaches such as PoseDiff, DUSt3R, MASt3R, FLARE, Fast3R, and CUT3R; (ii) monocular depth estimation, where we benchmark against DPT, P3Depth, SwinV2-L, AiT, VPD, ZoeDepth, and DepthAnything; and (iii) point-cloud reconstruction, where we consider pose-dependent methods such as MVDNet, GPMVS, Atlas, NeuralRecon, DG-Recon, and PanoRecon, together with pose-free approaches including RCVD, GCVD, COLMAP, FrozenRecon, and DUSt3R, ranging from traditional multi-view geometry to modern learning-based techniques. Notably, both depth and point-cloud evaluations are conducted in real scale.
Result & Analysis. Across the three benchmarks, Vid-LLM demonstrates strong reconstruction performance. On camera pose estimation (Tab. 6), it consistently surpasses both traditional optimization pipelines and advanced learning-based models on Co3Dv2 and RealEstate10K, confirming its robustness in preserving reliable multi-view consistency. On monocular depth estimation (Tab. 7), Vid-LLM achieves state-of-the-art results on NYU v2 at real scale, reaching an AbsRel of 0.025, RMSE of 0.109, log10 of 0.010, and a $\delta_{1}$ score of 0.987, representing substantial improvements over prior baselines. On point-cloud reconstruction (Tab. 8), Vid-LLM ranks as the best among pose-free methods on ScanNet, delivering an average improvement of 32% over the second-best approach, and even attains the highest Precision among all methods including pose-dependent pipelines, while maintaining a comparable global F-score. These results collectively indicate that Vid-LLM, despite operating without external poses, produces reliable real-scale depth and point clouds that are both clean and geometrically consistent.
@@ A.2 Supplementary Experiments on Reconstruction
Table 7: Depth estimation results on the NYU Depth v2 dataset.
@@
Table 8: Reconstruction results on the ScanNet dataset.
@@
Colmap (Schonberger & Frahm, 2016) $\times$ 0.142 0.367 0.119 0.267 0.178
@@
A.3 Implementation Details
In practice, we adopt DINOv2-L as the visual backbone, consisting of 24 Transformer layers with a hidden dimension of 1024. The projection MLPs use two fully connected layers with an expansion factor of 4 and GELU activation. For implementation, the Global–Frame Attention, camera head, and dense head in the reconstruction branch are adapted from VGGT (Wang et al., 2025a), while the construction of 3D patch representations in the reasoning branch is inspired by the approach used in LLaVA-3D (Zhu et al., 2025). All attention blocks are equipped with FlashAttention to reduce memory usage, and QK-Norm is applied prior to LayerNorm for stability. For data processing, we uniformly sample 32 frames per scene, resize the shorter side to 518, crop the resolution so that both height and width are multiples of 14, and apply standard augmentations including random cropping, color jittering, and horizontal flipping, ensuring augmentation independence across frames. Optimization is performed using AdamW with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and weight decay of $0.05$, with gradient clipping at $1.0$. We further enable mixed-precision training and activation checkpointing to reduce memory consumption.
A.4 The Use of Large Language Models (LLMs) Statement

### v2

@@ Appendix A Appendix
  • qualitative results and challenging cases of Vid-LLM in Appendix A.1;
@@
A.1 Qualitative Results and Challenging Cases
@@
Figure 8: Challenging Cases. We present one challenging case for each 3D VL task—3D question answering (a), 3D visual grounding (b), and 3D dense captioning (c). (a) Due to limited camera viewpoints in the input video, one chair is only partially observed and consequently missing from the reconstruction, leading to an undercounted result. (b) The highly reflective tabletop results in unstable depth estimates and incomplete reconstruction, resulting in a predicted 3D box that is smaller than the ground truth. (c) The top surface of the cabinet is not well reconstructed because it appears only briefly and mostly from oblique angles, yet Vid-LLM nonetheless produces an accurate caption by leveraging rich 2D semantic cues.
Qualitative Results. Fig. 7 presents reconstruction results on the ScanNet dataset, together with three representative 3D vision–language (3D VL) tasks. Specifically, it illustrates (1) 3D Question Answering, where the model answers queries on object counts, locations, and attributes from reconstructed scenes; (2) 3D Dense Captioning, which provides detailed semantic descriptions of rooms and key objects; and (3) 3D Visual Grounding, where the model localizes target objects in 3D space according to textual instructions. These qualitative results demonstrate the capability of Vid-LLM in scene reconstruction and 3D VL reasoning from video inputs.
Challenging Cases. Fig. 8 presents three challenging cases for Vid-LLM on 3D VL tasks. These scenes reflect typical failure modes of monocular video reconstruction, where limited camera viewpoints, reflective surfaces, or oblique viewing angles lead to incomplete geometry. As shown in examples (a) and (b), tasks that rely on explicit 3D structural information, such as instance counting or estimating 3D object extents for grounding, are directly affected by incomplete or unreliable geometry. In contrast, example (c) shows that Vid-LLM remains effective when sufficient 2D cues are present, due to the Cross-Task Adapter and bridge-token mechanism, which jointly align semantic and geometric information. These observations suggest that improving the reconstruction branch to enhance geometric fidelity could further strengthen the overall 3D VL performance, indicating a promising direction for future work.
@@
Table 7: Depth estimation results on the NYU Depth v2 dataset. Model outputs are in relative scale; we align them to the ground-truth metric scale prior to evaluation.
@@
Table 8: Reconstruction results on the ScanNet dataset. Model outputs are in relative scale; we align them to the ground-truth metric scale prior to evaluation.
@@
Colmap  (Schonberger and Frahm, 2016) $\times$ 0.142 0.367 0.119 0.267 0.178
@@
A.2 Supplementary Experiments on Reconstruction
Overview. Since the reconstruction branch not only provides 3D cues for the 3D-MLLM but also explicitly produces point clouds that can serve downstream applications, it is important to assess its individual performance. To evaluate the 3D modeling capability of Vid-LLM, we assess three tasks: camera pose estimation, depth prediction, and point-cloud reconstruction. Depth prediction is crucial to real-scale reconstruction, and together with poses it determines the quality of the recovered point clouds. Following standard protocols, we evaluate poses on Co3Dv2 (Reizenstein et al., 2021) and RealEstate10K (Zhou et al., 2018), depth on the indoor NYU Depth v2 dataset (Silberman et al., 2012), and point-cloud reconstruction on ScanNet (Dai et al., 2017), ensuring consistency with the 3D vision–language benchmarks introduced earlier. Depth and reconstruction evaluations require the model to predict real-scale geometric outputs, while camera pose estimation is scale-invariant.
Baseline. For 3D reconstruction, we evaluate three subtasks: (i) camera pose estimation, where we compare against geometry-based optimization methods including COLMAP+SG and PixSfM, as well as state-of-the-art learning-based approaches such as PoseDiff, DUSt3R, MASt3R, FLARE, Fast3R, and CUT3R; (ii) monocular depth estimation, where we benchmark against DPT, P3Depth, SwinV2-L, AiT, VPD, ZoeDepth, and DepthAnything; and (iii) point-cloud reconstruction, where we consider pose-dependent methods such as MVDNet, GPMVS, Atlas, NeuralRecon, DG-Recon, and PanoRecon, together with pose-free approaches including RCVD, GCVD, COLMAP, FrozenRecon, and DUSt3R, ranging from traditional multi-view geometry to modern learning-based techniques. We further benchmark against VGGT, which serves as both a high-performing reconstruction model and as the reconstruction teacher in our framework.
Result & Analysis. Across the three benchmarks, Vid-LLM achieves reconstruction quality comparable to VGGT while operating in an end-to-end real-scale setting. On camera pose estimation (Tab. 6), Vid-LLM attains results close to VGGT on both Co3Dv2 and RealEstate10K and slightly outperforms it on RTA@15 for Co3Dv2. On NYU Depth v2 and ScanNet (Tab. 7 and Tab. 8), VGGT requires an explicit scale-alignment step to match the metric ground truth, whereas Vid-LLM predicts real-scale depth and point clouds in an end-to-end manner. In this setting, Vid-LLM achieves lower log10 error on NYU and higher F-score and accuracy on ScanNet than VGGT, indicating that its real-scale geometry remains competitive against the teacher model. Vid-LLM also surpasses all other models that predict real-scale outputs. Overall, these results confirm that its reconstruction module delivers consistent and reliable geometric performance.
A.3 The Use of Large Language Models (LLMs) Statement